OpenStudy (penguin1234):

Which situation is linear? A: A sample of bacteria is doubling every day B: the number of cars each hour on a road for a week C: the cost of a movie ticket increases five percent each year D: the population of a town in increasing by ten people each year

OpenStudy (penguin1234):

What is the answer

OpenStudy (anonymous):

A is going to be a exponential regression. B is just random dots (data points).

OpenStudy (penguin1234):

Sure

OpenStudy (penguin1234):

So D would be the answer

OpenStudy (anonymous):

C is also an exponential regression because it deals with percentages. and D is the correct answer because your equation will be y=10x with x=the year

OpenStudy (penguin1234):

Thank You

OpenStudy (the_fizicx99):

Yes. See, A is exponential because it's "doubling" If it started of with, say 2, when 2 * 2 = 4, * 2 = 8, * 2 = 16 and so on. The data would be{2,4,8,16...} Linear is a \(\ \sf constant\) rate of change therefore it's either being added or subtract the same amount. B would be random points because it would vary. C is also exponential. D is increasing by 10 each year, so it's a constant +10 each time so that is linear. You could set it up as, y = 10x, if you had a starting value that would be the b, in y = mx + b, b is the y-intercept. (0,?)

OpenStudy (anonymous):

You're welcome (: